# Bier spheres and Toric Topology

### v1

@@ 3. General case: Buchstaber number
In this section we turn to the discussion of combinatorial properties of Bier spheres in arbitrary dimensions in their relation with topological properties of certain toric spaces associated to the Bier spheres by means of toric topology.
The key notion of study in toric topology is that of a polyhedral product. A particularly important case of a polyhedral product for us here is the following object.
@@
Here we denote by $\mathbb{D}^{2}=\{z\in\mathbb{C}\,|\,|z|\leqslant 1\}$ the unit disc in the complex plane and by $\mathbb{S}^{1}$ – its boundary. Note that $\mathcal{Z}_{\mathcal{K}}$ can be naturally viewed as an $(m+n)$-dimensional cellular subspace in the unitary polydisc $(\mathbb{D}^{2})^{m}$ in the complex Euclidean space $\mathbb{C}^{m}$.
@@
Set $\mathcal{K}_{1}:=\mathcal{K}$ and $\mathcal{K}_{2}:=\widehat{\mathcal{K}}$, its Alexander dual. Then their deleted join equals $\mathrm{Bier}(\mathcal{K})$, the Bier sphere of $\mathcal{K}$, which is a PL-sphere of dimension $m-2$. Then, by the previous lemma, $\mathrm{Bier}(\mathcal{K})$ has a maximal possible (real) Buchstaber number. Q.E.D. ∎
@@ 4. Examples, applications and open problems
In this section we consider some examples of our main results application and formulate some open problems. Let us begin with the following example.
@@
Applying Theorem 2.18, we get $13$ characteristic maps $\Lambda_{\mathcal{P}_{i}}$, quasitoric manifolds $\mathcal{M}_{i}$ and small covers $\overline{\mathcal{M}_{i}}$. Note that for $i\neq 4,6$ they are canonical for the canonical Delzant realizations of $\mathcal{P}_{i}$, described above. In the case $i=4,6$ we take the realizations from 2.17 and 2.15, respectively.
@@
We compute the $f$-vectors of all the 13 Bier spheres $\mathcal{S}_{i}$, see Appendix A. Using them, one can immediately restore the corresponding $h$-vectors and show that the $6$-dimensional quasitoric manifolds $\mathcal{M}_{i}$ for $1\leqslant i\leqslant 13$ are nonsingular projective toric varieties having integer homology groups, shown in the table below:
@@
Thus, the mod 2 reduction of the characteristic matrices collected in Appendix B shows that among the $3$-dimensional small covers $\overline{\mathcal{M}_{i}}$, the 1st, 6th, 9th, 12th and 13th are orientable and the rest are not.
@@ 6. Appendix B
    $\Lambda_{\mathcal{P}_{4}}=\begin{pmatrix}1&1&0&-1&0&0&0&1\\ 0&0&0&0&0&-1&1&1\\ 0&1&1&0&-1&0&0&0\end{pmatrix}\in\mathrm{Mat}_{3\times 8}(\mathbb{Z});$
@@
    $\Lambda_{\mathcal{P}_{11}}=\begin{pmatrix}1&0&0&-1&1&-1\\ 0&1&0&-1&1&0\\ 0&0&1&-1&1&-1\end{pmatrix}\in\text{Mat}_{3\times 6}(\mathbb{Z});$
@@ Acknowledgements
The authors are grateful to Anton Ayzenberg, Taras Panov, and Rade Živaljević for numerous fruitful discussions, valuable comments and suggestions. The authors also express their gratitude to the anonymous referee for the comments and suggestions that made it possible to improve the presentation.
Limonchenko was supported by the Serbian Ministry of Science, Innovations and Technological Development through the Mathematical Institute of the Serbian Academy of Sciences and Arts. Sergeev was supported within the framework of the HSE University Basic Research Program.

### v2

@@ 3. General case: Buchstaber number
In this section we turn to a discussion of the combinatorial properties of Bier spheres related to the topological properties of their polyhedral products studied in the framework of toric topology.
The key notion of study in toric topology [4, 5] is that of a polyhedral product. A particularly important case of a polyhedral product for us here is the following object.
@@
Here we denote by $\mathbb{D}^{2}=\{z\in\mathbb{C}\,|\,|z|\leqslant 1\}$ the unit disc in the complex plane and by $\mathbb{S}^{1}$ its boundary. Note that $\mathcal{Z}_{\mathcal{K}}$ can be naturally viewed as an $(m+n)$-dimensional cellular subspace in the unitary polydisc $(\mathbb{D}^{2})^{m}$ in the complex Euclidean space $\mathbb{C}^{m}$.
@@
Set $\mathcal{K}_{1}:=\mathcal{K}$ and $\mathcal{K}_{2}:=\widehat{\mathcal{K}}$, its Alexander dual. Then their deleted join equals $\mathrm{Bier}(\mathcal{K})$, the Bier sphere of $\mathcal{K}$, which is a PL-sphere of dimension $m-2$. Then, by the previous lemma, $\mathrm{Bier}(\mathcal{K})$ has the maximal possible (real) Buchstaber number. Q.E.D. ∎
@@ 4. Examples, applications and open problems
In this section we consider some examples of applications of our main results and formulate some open problems. Let us begin with the following example.
@@
Applying Theorem 2.16, we get $13$ characteristic maps $\Lambda_{\mathcal{P}_{i}}$, quasitoric manifolds $\mathcal{M}_{i}$ and small covers $\overline{\mathcal{M}_{i}}$. Note that for $i\neq 6$ they are canonical for the canonical Delzant realizations of $\mathcal{P}_{i}$, described above. In the case of $i=6$ we take the realization from Proposition 2.15.
@@
We compute the $f$-vectors of all the 13 Bier spheres $\mathcal{S}_{i}$, see Appendix A. Using them, one can immediately restore the corresponding $h$-vectors and show that the six-dimensional quasitoric manifolds $\mathcal{M}_{i}$ for $1\leqslant i\leqslant 13$ are nonsingular projective toric varieties with the integer homology groups shown in Table 3.
@@
Table 3. Integral homology groups of the quasitoric manifolds $\mathcal{M}_{i}$
@@
Thus, the mod 2 reduction of the characteristic matrices collected in Appendix B shows that among the $3$-dimensional small covers $\overline{\mathcal{M}_{i}}$, the 1st, 6th, 9th, 11th, 12th and 13th are orientable and the rest are not.
@@ 6. Appendix B
    $\Lambda_{\mathcal{P}_{4}}=\begin{pmatrix}1&0&0&-1&1&1&-1&0\\ 0&1&0&-1&1&1&-1&-1\\ 0&0&1&-1&0&1&0&0\end{pmatrix}\in\mathrm{Mat}_{3\times 8}(\mathbb{Z});$
@@
    $\Lambda_{\mathcal{P}_{11}}=\begin{pmatrix}1&0&0&-1&1&-1\\ 0&1&0&-1&1&0\\ 0&0&1&-1&1&0\end{pmatrix}\in\text{Mat}_{3\times 6}(\mathbb{Z});$
@@ Acknowledgements
The authors are grateful to Anton Ayzenberg, Taras Panov, and Rade Živaljević for numerous fruitful discussions, valuable comments and suggestions. The authors also express their gratitude to Younghan Yoon and the anonymous referee for the comments and suggestions that made it possible to improve the presentation.
Limonchenko was supported by the Serbian Ministry of Science, Innovations and Technological Development through the Mathematical Institute of the Serbian Academy of Sciences and Arts. Sergeev‘s work was carried out within the project ‘‘Mirror Laboratories’’ of HSE University, Russian Federation.